OpenStudy (anonymous):

lim sin2x/sinx x--->0

OpenStudy (campbell_st):

well us the substitution sin(2x) = 2sin(x)cos(x) then see what happens...

OpenStudy (anonymous):

that's an identity ? never knew that.

OpenStudy (campbell_st):

ok.... its a double angle identity

OpenStudy (anonymous):

okay i got it is 2, thanks! ;D

OpenStudy (campbell_st):

is the question 1. \[\frac{\sin(2x)}{\sin(x)}\] or 2. \[\frac{\sin^2(x)}{\sin(x)}\]

OpenStudy (anonymous):

1 lol

OpenStudy (anonymous):

i know how to make an exponent lol

OpenStudy (campbell_st):

ok...

OpenStudy (anonymous):

thank you though you helped me

OpenStudy (anonymous):

how about sin4x/sin3x ?is there an identity for that lol